OpenStudy (anonymous):

PLEASE HELP!!! ILL GIVE MEDAL!!!!! Which expression is equivalent to this one? u · 5 + u · v A. u(5 + v) B. 5(u + v) C. v(5 + u) D. u · (5 + u) · v

OpenStudy (anonymous):

the answer would be A cause they use the distributive property which another version of the example

OpenStudy (elonasushchik):

It's B .

OpenStudy (anonymous):

oh my fault I thought a was b

OpenStudy (elonasushchik):

Medal?!

OpenStudy (anonymous):

lol ok can u help me with 3 more????

OpenStudy (anonymous):

yes the correct answer is B

OpenStudy (anonymous):

:)

OpenStudy (elonasushchik):

Sure @Kayla369

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

What is the solution to the equation? 4(3 + 7) = 4 · x + 4 · 7 A. x = 3 B. x = 4 C. x = 7 D. x = 10

OpenStudy (anonymous):

What is the solution to the equation? 4(3 plus 7) = 4 times x plus 4 times 7 A. x = 3 B. x = 4 C. x = 7 D. x = 10

OpenStudy (elonasushchik):

@nechirwan umm. Can u help..

OpenStudy (anonymous):

xD haha ya it confusing lol

OpenStudy (anonymous):

1 is A //2 is C I believe

OpenStudy (anonymous):

1 is A?????? i thought it was B

OpenStudy (anonymous):

ok well i have another one

OpenStudy (anonymous):

Use the distributive property to rewrite the expression and solve. 9 72 = 9 times ???? + 9 times2 = ????

OpenStudy (anonymous):

o ya n 9 times 71 also haha

OpenStudy (anonymous):

72*

OpenStudy (anonymous):

so basically 9 times 72 = 9 times ????? + 9 times 2 = ?????